For question 2, please draw the graph. 4) Suppose Microsoft computers has a cost curve equal to C = 150Q and faces a demand curve equal to p = 550 – 22. The Bay Are
Answer #1

Com buty has a cost cure I - suppose equal to Microsoft c = 150 Q and faces a demand cuuc equal to þ- 550-20 02 Now, To tal rNow, Darand cuand up-550 2Q Quant apie Q = 275 When P=0 = OG Q20 > P-550 Also MRS 550-4Q When Q=0 => MR=55000 - when MR 0 ->l et the regulaston sit the price eques to the - marginal cost, - then we will have !! P- ML = 150 - s pz 150 : 2 and optimali Nows Pio bit maximising for Probit maximising condition will beirt Cond MRE MC : le gestellt 550-4Q = 160 21 3):4Q390 _Q_ 9